Circular No. 8826

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER AND SATURN Further to IAUC 8502 and 8730, the IAU Working Group for Planetary

System Nomenclature (WGPSN) has approved the following new designations and names of satellites of Jupiter and Saturn:

Jupiter XLIX  Kore  = S/2003 J 14
Saturn XXXVI Aegir  = S/2004 S 10
Saturn XXXVII Bebhionn = S/2004 S 11
Saturn XXXVIII Bergelmir = S/2004 S 15
Saturn XXXIX Bestla  = S/2004 S 18
Saturn XL  Farbauti = S/2004 S 9
Saturn XLI  Fenrir  = S/2004 S 16
Saturn XLII  Fornjot  = S/2004 S 8
Saturn XLIII  Hati  = S/2004 S 14
Saturn XLIV  Hyrokkin = S/2004 S 19
Saturn XLV  Kari  = S/2006 S 2
Saturn XLVI  Loge  = S/2006 S 5
Saturn XLVII  Skoll  = S/2006 S 8
Saturn XLVIII  Surtur  = S/2006 S 7

RING OF URANUS The IAU WGPSN has recently designated the Uranian ring 1986 U 2R

(cf. IAUC 8373; Stone and Miner 1986, Science 233, 39) the “ζ ring”.

2006 VV_{2}

L. A. M. Benner, S. J. Ostro, and J. D. Giorgini, Jet Propulsion Laboratory (JPL), California Institute of Technology (CIT); M. W. Busch, CIT; R. Rose, M. A. Slade, and R. F. Jurgens, JPL/CIT; M. C. Nolan and A. A. Hine, Arecibo Observatory; G. J. Black, University of Virginia; and L. M. Carter, Smithsonian Institution, report that Goldstone (8560 MHz, 3.5 cm) and Arecibo (2380 MHz, 12.6 cm) radar observations during Mar. 27-Apr. 3 reveal that the minor planet 2006 VV_{2} (cf. MPEC 2006-V47, 2006-W87), which passed only 0.0226 AU from the earth around Mar. 31.25 UT (with computed V = 10.0) and has H = 16.7 (MPO 118180), is a binary system. Preliminary estimates for the component diameters are 1.8 and > 0.3 km. The maximum orbital distance between the components is at least 1.5 km.

2007 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green